# Probabilistic Soundness Guarantees in LLM Reasoning Chains

### v1

@@ Abstract
In reasoning chains generated by large language models (LLMs), initial errors often propagate and undermine the reliability of the final conclusion. Current LLM-based error detection methods often fail to detect propagated errors because they do not properly account for how earlier errors might corrupt judgments of downstream reasoning. To better detect such propagated errors, we introduce Autoregressive Reasoning Entailment Stability (ARES), a novel probabilistic framework that prevents error propagation by judging each claim based only on previously-assessed sound premises. This inductive method yields a nuanced score for each step and provides certified statistical guarantees of its soundness, rather than a brittle binary label. ARES achieves state-of-the-art performance across four benchmarks (72.1% Macro-F1, +8.2 points) and demonstrates superior robustness on very long synthetic reasoning chains, where it excels at detecting propagated errors (90.3% F1, +27.6 points). 111Correspondence to weiqiuy@seas.upenn.edu. Code is available at https://github.com/fallcat/ares.
@@
Weiqiu You Anton Xue Shreya Havaldar Delip Rao Helen Jin Chris Callison-Burch Eric Wong University of Pennsylvania
@@ 1 Introduction
Large Language Models (LLMs) frequently generate chain-of-thought reasoning processes alongside their final outputs, yet these often contain errors (Huang et al., 2025; Lyu et al., 2024). This critically compromises the trustworthiness and reliability of LLM-generated content, consequently diminishing user confidence and impeding the broader adoption of LLMs in high-stakes applications (Agarwal et al., 2024; Chen and Mueller, 2023).
@@
As illustrated in Figure 1, one type of error is an ungrounded step—a step that is incorrect with respect to the given context. For example, the model might incorrectly copy a 2/5 in the context to be 3/5. Another common error is an invalid derivation, such as a logical misstep or miscalculation—for example, deriving from (Lee and Hockenmaier, 2025). A third type of error involves error propagation: even if the logic is valid, an incorrect starting assumption can lead to a wrong conclusion. For instance, using the incorrect claim to derive is logically valid but the derived claim is incorrect due to the initial error (Tyagi et al., 2024).
@@
Current error detection methods typically aim to identify all errors at once. For example, LLM judges are prompted to evaluate the entire chain and assess each step for correctness (Tyagi et al., 2024; He et al., 2025). Similarly, Process Reward Models (PRMs) are language models trained with step-level classification heads on this same objective (Lightman et al., 2023).
@@
However, existing error detection methods often fall short. Specifically, they are often distracted by the presence of propagated errors (He et al., 2025; Turpin et al., 2023; Dhuliawala et al., 2023). In the example from Figure 1, if steps 3, 4, and 5 are evaluated together, an LLM may incorrectly mark step 5 as sound by incorrectly relying on step 4, which is invalid. This highlights the need for robust methods that can assess the soundness of each step without being adversely distracted by prior errors.
@@
To address this issue, we draw inspiration from human reasoning. Humans typically review claims sequentially, and disregard previously unsound statements when evaluating subsequent ones (Johnson-Laird, 2010; Mukherjee et al., 2025). In contrast, LLMs struggle to ignore prior errors, which causes naive detection methods to fail at simultaneously identifying and localizing all errors in a reasoning chain (Wu et al., 2024; Song and Tavanapong, 2024). To overcome this limitation, we develop Autoregressive Reasoning Entailment Stability (ARES), a probabilistic framework that evaluates the soundness of each reasoning step based on its expected entailment probability, conditioned only on previously-occurring, sound claims (Figure 2). We iteratively evaluate each claim as follows: entailed claims are retained as premises for subsequent steps, while non-entailed claims are discarded. For uncertain claims, retention is probabilistic based on the entailment model. This adaptation not only improves error detection but also enables us to give certified guarantees on the robustness of a reasoning chain.
@@ 2 Soundness in Reasoning Chains
We aim to identify and certify errors within LLM-generated chain-of-thought (CoT) reasoning. To this end, this section formalizes reasoning chains in terms of their constituent claims (Section 2.1), introduces the concept of probabilistic entailment between these claims (Section 2.2), and defines a notion of soundness that incorporates internal groundedness, validity, and the entailment of a final hypothesis (Section 2.3).
@@ 2.3 Reasoning Chains and Soundness
This partition is methodologically crucial. Base claims () serve as the foundational premises for a given reasoning task; their factual accuracy is considered out of scope for the present analysis and is assumed to be handled by external mechanisms. We focus on assessing whether each derived claim ( for ) is soundly inferred from the set of preceding statements. We begin by defining a deterministic (i.e., “hard”) version of soundness, where the entailment model is assumed to be binary-valued.
@@
Consider a reasoning chain with base claims and derived claims . Then, this reasoning chain is hard-sound with respect to the deterministic entailment model if
for all derived claims indexed by .
The concept of hard soundness provides a precise, albeit strict, benchmark for evaluating the logical integrity of a reasoning chain because it requires every derived claim to be perfectly entailed by its predecessors. However, LLM-generated reasoning chains often deviate from this ideal. Therefore, while hard soundness serves as an important theoretical standard of correctness, it cannot give nuanced measures of error, particularly in long reasoning chains. This necessitates more flexible methods for measuring claim soundness even in the presence of errors, which we address next.
@@ 3 Soundness Checks via Autoregressive Reasoning Entailment Stability
for reasoning steps . We aim to quantify the reliability of this process using a sequence of entailment stability scores: , where each denotes how reliably the -th derived claim () is entailed with respect to its preceding claims . The connection between entailment and error detection is straightforward: a claim is likely to be erroneous if is low.
While the notion of reliability is general, a rigorous definition of each is challenging due to probabilistic uncertainty and unreliable preceding claims in the reasoning chain. Critically, deterministic hard soundness (Definition 2.1) cannot account for premises with uncertainty. Autoregressive Reasoning Entailment Stability addresses this: Section 3.1 motivates probabilistic entailment using insights from human psychology, LLM empirics, and mathematical logic. Subsequently, Section 3.2 formalizes our approach, defines Autoregressive Reasoning Entailment Stability, and details its efficient Monte Carlo estimation.
@@ 3.1 Entailment with Probabilistic Premises
The key challenge lies in accurately assessing entailment when premises are probabilistically uncertain. To address this, our main insight is to calculate an overall likelihood by averaging across various probable combinations of that uncertain information.
@@ 3.2 Autoregressive Reasoning Entailment Stability with Efficient Sampling
In the previous section, we established a method for calculating the entailment of a single hypothesis based on a set of premises that might be uncertain (Equation 3). Now, we will extend this concept to evaluate an entire LLM-generated reasoning chain, which consists of multiple, sequential steps. Our goal is to compute a sequence of entailment stability scores, denoted as , where each score quantifies the reliability of the -th derived claim, .
The core challenge remains the same: how do we reliably judge a claim when the preceding claims it relies on are themselves not entirely trustworthy? Our approach, Autoregressive Reasoning Entailment Stability (ARES), solves this by autoregressively assessing each claim while accounting for the soundness of previous claims. In particular, when we evaluate the -th derived claim, we consider all possible combinations of soundness for the preceding claims. The stability score, , is then the expected entailment of the current claim, averaged across all sound combinations.
@@
Inductive Case (): For any derived claim after the first, the probability of a specific premise combination depends on two factors: the probability of the preceding combination () and the entailment probability of the new claim given that preceding combination. In other words, a claim is added to our set of “sound” premises for the next step based on how strongly the currently accepted set entails it:
(5) where denotes the subset of claims indexed by .
@@
However, directly computing is highly inefficient, as it requires summing over possible premise combinations. Instead, we estimated it by sampling the premise combinations:
where let be i.i.d. sampled according to Algorithm 1 and Figure 3. Additionally, note that converges rapidly to as the number of samples grows, allowing us to obtain a rigorous statistical guarantee on our stability scores as a function of the number of samples.
@@
Let for any and . For any entailment model and reasoning chain with base claims and derived claims. Then, define the stability estimators for all derived claims as in Equation 7. Then, with probability at least , we have accuracy for all .
@@
Recall the connection between entailment stability and error detection: the lower a claim’s entailment stability, the greater its error. Consider a simple thresholding mechanism: if some falls below a prescribed error threshold, then we mark the derived claim as erroneous. In the following, we demonstrate the empirical effectiveness of this procedure.
@@ 4 Evaluating ARES for Estimating Probabilistic Soundness
For models, we include both proprietary LLM GPT-4o-mini (OpenAI, 2024) and open source LLM Qwen3-4B (Yang et al., 2025), as well as Process Reward Model Qwen2.5-Math-PRM-7B (Zhang et al., 2025). We evaluate on four datasets: PRMBench (Song et al., 2025), DeltaBench (He et al., 2025), ClaimTrees (our synthetic data), and CaptainCookRecipes (graph-based recipe dataset derived from CaptainCook4D (Peddi et al., 2024)). We evaluate using Macro-recall, Macro-precision and Macro-F1 following the literature (He et al., 2025). To compute the error threshold for the entailment scores, we first sweep over all the values that occur in the training split and select the one that maximizes Macro-F1. We repeat this process in a 5-fold cross-validation and report the average and standard deviation. Additional details and analyses can be found in Appendix C.
@@ 4.1 RQ1: Does ARES work better than baseline methods on Benchmarks?
We measure ARES’s ability to identify errors in natural reasoning chains using PRMBench and DeltaBench. With GPT-4o-mini backbone entailment model, we find that ARES achieves the best Macro-F1 scores on both datasets, shown in Table 1. LLM-Judge performs poorly on DeltaBench while Entail-Base underperforms on PRMBench. DeltaBench’s long reasoning chains appear to confuse LLM-Judge when making holistic judgments. For Qwen3-4B, Entail-Base performs slightly better, while all other methods lag behind. Our inspection reveals that Qwen3-4B-based entailment models frequently classify next claims as entailed, suggesting limited capability for judging complex reasoning. Additional experiments in Section C.8 show that ARES, built on a PRM backbone, achieves further improvements over it.
@@ 4.2 RQ2: In what setting does ARES identify more errors than baselines?
CaptainCookRecipes: A dataset adapted from the cooking task graphs in CaptainCook4D (Peddi et al., 2024), which rely on sequential rules.
We designed these datasets specifically to test the core reasoning capabilities of each method. To control for confounding variables like world knowledge and greedy reasoning , ClaimTrees uses abstract symbols and shuffled rules to mitigate ordering bias. In both datasets, we represent the underlying rules (e.g., logical rules or recipe actions) as base claims. We then intentionally remove a key base claim—such as a rule in a proof or an ingredient in a recipe—to create unsound derivations where we can track exactly how the initial error propagates through the reasoning chain. Further details on these datasets are available in Appendix C.4.
Experiments on these controlled datasets confirm that ARES excels at identifying propagated errors, especially in long chains. As demonstrated in Figure 4, ARES maintains a high Macro-F1 score even as chains become very long, whereas the performance of all baseline methods deteriorates sharply after just a few steps. For example, ARES sustains a Macro-F1 score of at least 89% on chains up to 50 steps long, while other methods fall into the 30–40% range. The results in Table 1 further highlight this robust performance across our synthetic datasets. We further discuss in Appendix B that only ARES satisfies all important desiderata for error detection while other methods fail to.
4.3 RQ3: Is ARES computationally efficient?
While ARES samples multiple combinations of previous claims to check soundness, it is implemented efficiently to avoid redundant LLM calls for the same premise-hypothesis pairs. Figure 5 shows the theoretical versus actual samples used for each derived claim and complete reasoning chain, respectively. For ClaimTrees, the average total samples per example increases with chain length. With shorter chains (ClaimTrees-5), we achieve extreme efficiency at only 0.03x of theoretical samples needed. DeltaBench uses the most samples but still achieves 0.31x of theoretical samples needed. DeltaBench needing more samples indicates greater uncertainty in the entailment model’s outputted probability for this dataset. In ideal cases where the entailment model outputs only 1 or 0 for every derived claim, we need just one sample per claim since we do not need to compute entailment for any claim combinations in the other case.
@@ 4.3 RQ3: Is ARES computationally efficient?
To have a more direct analysis of efficiency, we conduct another analysis on the performance vs. sample size trade-off across all datasets on GPT-4o-mini Table 2. We find that in practice ARES’s performance is stable even with fewer samples, indicating potential further computational savings. On synthetic benchmarks ClaimTrees and CaptainCookRecipes, there is no significant difference for to , while more difference for PRMBench and DeltaBench.

### v2

@@ Abstract
In reasoning chains generated by large language models (LLMs), initial errors often propagate and undermine the reliability of the final conclusion. Current LLM-based error detection methods often fail to detect propagated errors because earlier errors can corrupt judgments of downstream reasoning. To better detect such errors, we introduce Autoregressive Reasoning Entailment Stability (ARES), a probabilistic framework that evaluates each reasoning step based solely on previously-verified premises. This inductive method yields a nuanced score for each step and provides certified statistical guarantees of its soundness, rather than a brittle binary label. ARES achieves state-of-the-art performance across four benchmarks (72.1% Macro-F1, +8.2 points) and demonstrates superior robustness on very long synthetic reasoning chains, where it excels at detecting propagated errors (90.3% F1, +27.6 points). 111Correspondence to weiqiuy@seas.upenn.edu. Code is available at https://github.com/fallcat/ares.
@@
Weiqiu You† Anton Xue⋆ Shreya Havaldar† Delip Rao† Helen Jin† Chris Callison-Burch† Eric Wong† †University of Pennsylvania ⋆University of Texas at Austin
@@ 1 Introduction
Large Language Models (LLMs) are taking on increasingly sophisticated reasoning tasks in critical fields like medicine and scientific discovery. Yet, a fundamental challenge persists: the chain-of-thought processes that lead to their outputs are frequently flawed with errors (Huang et al., 2025; Lyu et al., 2024). This critically compromises the reliability of LLM-generated content, diminishing user confidence and impeding the broader adoption of LLMs in high-stakes applications (Agarwal et al., 2024; Chen and Mueller, 2024).
@@
As illustrated in Figure˜1, one type of error is an ungrounded step—a step that is incorrect with respect to the given context. For example, the model might incorrectly copy a 2/5 in the context to be 3/5. Another common error is an invalid derivation—for example, deriving from —which is a logical misstep or miscalculation (Lee and Hockenmaier, 2025). A third type of error involves error propagation: even if the logic is valid, an incorrect starting assumption can lead to a wrong conclusion. For instance, using the incorrect claim to derive is logically valid but the derived claim is incorrect due to the initial error (Tyagi et al., 2024).
@@
Current error detection methods typically aim to identify all errors at once. For example, LLM judges are prompted to evaluate the entire chain and assess each step for correctness (Tyagi et al., 2024; He et al., 2025). Similarly, Process Reward Models (PRMs) are language models trained with step-level classification heads on this same objective (Lightman et al., 2024).
@@
However, existing error detection methods often fall short. Specifically, they are often distracted by the presence of propagated errors (He et al., 2025; Turpin et al., 2023; Dhuliawala et al., 2024). In the example from Figure 1, if steps 3, 4, and 5 are evaluated together, an LLM may incorrectly mark step 5 as sound by incorrectly relying on step 4, which is invalid. This highlights the need for robust methods that can assess the soundness of each step without being adversely distracted by prior errors.
@@
To address this issue, we draw inspiration from human reasoning. Humans typically review claims sequentially, and disregard previously unsound statements when evaluating subsequent ones (Johnson-Laird, 2010; Mukherjee et al., 2025). In contrast, LLMs struggle to ignore prior errors, which causes naive detection methods to fail at simultaneously identifying and localizing all errors in a reasoning chain (Wu et al., 2024a; Song and Tavanapong, 2024). To overcome this limitation, we develop Autoregressive Reasoning Entailment Stability (ARES), a probabilistic framework that evaluates the soundness of each reasoning step based on its expected entailment probability, conditioned only on previously-occurring, sound claims (Figure 2). We iteratively evaluate each claim as follows: entailed claims are retained as premises for subsequent steps, while non-entailed claims are discarded. For uncertain claims, retention is probabilistic based on the entailment model. This adaptation not only improves error detection but also enables us to give certified guarantees on the robustness of a reasoning chain.
@@ 2 Soundness in Reasoning Chains
We aim to identify and certify errors within LLM-generated chain-of-thought (CoT) reasoning. To this end, this section formalizes reasoning chains in terms of their constituent claims (Section˜2.1), introduces the concept of probabilistic entailment between these claims (Section˜2.2), and defines a notion of soundness that incorporates internal groundedness, validity, and the entailment of a final hypothesis (Section˜2.3).
@@ 2.3 Reasoning Chains and Soundness
This partition is methodologically crucial. Base claims () serve as the foundational premises for a given reasoning task; their factual accuracy is given and assumed to be validated by external mechanisms. Instead, we focus on assessing whether each derived claim ( for ) is soundly inferred from the set of preceding statements. To begin, we define a deterministic (i.e., “hard”) version of soundness, where all derived claims are entailed with certainty.
@@
A reasoning chain is hard-sound with respect to the entailment model if for all derived claims, we have
The concept of hard soundness provides a precise, albeit strict, benchmark for evaluating the logical integrity of a reasoning chain: it requires every derived claim to be perfectly entailed by its predecessors. However, LLM-generated reasoning chains often deviate from this ideal. Therefore, while hard soundness serves as an important theoretical standard of correctness, it cannot give nuanced measures of error, particularly in long reasoning chains. This necessitates more flexible methods for measuring claim soundness even in the presence of errors, which we address next.
@@ 3 Soundness Checks via Autoregressive Reasoning Entailment Stability
for reasoning steps . We aim to quantify the reliability of this process using a sequence of entailment stability scores: , where each denotes how reliably the -th derived claim is entailed with respect to its preceding claims . The connection between entailment and error detection is straightforward: if is small, then is likely an error.
However, a well-principled and computationally tractable formulation of is non-trivial when entailment is probabilistic. Critically, hard soundness is incompatible with non-binary outputs, and it is not immediately clear how uncertain premises should be evaluated. ARES addresses this: Section˜3.1 motivates probabilistic entailment using insights from human psychology, LLM empirics, and mathematical logic. Subsequently, Section˜3.2 formalizes our approach, defines ARES, and details its efficient Monte Carlo estimation.
@@ 3.1 Entailment with Probabilistic Premises
The key challenge lies in accurately assessing entailment when premises are probabilistically uncertain. Our main insight is to calculate an overall likelihood by averaging across various probable combinations of that uncertain information.
@@ 3.2 Autoregressive Reasoning Entailment Stability with Efficient Sampling
We previously established a method for calculating the entailment of a single hypothesis based on a set of premises that might be uncertain (Equation˜3). Now, we will extend this concept to evaluate an entire LLM-generated reasoning chain, which consists of multiple steps. Our goal is to compute a sequence of entailment stability scores, denoted as , where each score quantifies the reliability of the -th derived claim, .
The core challenge remains: how to reliably judge a claim when the preceding claims it relies on are themselves not entirely trustworthy? Our approach, ARES, solves this by autoregressively assessing each claim while accounting for the soundness of previous claims. In particular, when we evaluate the -th derived claim, we consider all possible combinations of soundness for the preceding claims. The stability score, , is then the expected entailment of the current claim, averaged across all sound combinations.
@@
Inductive Case (): For subsequent claims, the probability of a specific premise combination depends on two factors: the probability of the preceding combination () and the entailment probability of the new claim given that preceding combination. That is, a claim is added to our set of “sound” premises based on how strongly the current set entails it, where let
(5) where are the claims indexed by .
@@
However, directly computing is highly inefficient, as it requires summing over possible combinations of premise entailment. Instead, we estimated it by sampling the premise combinations:
where let be i.i.d. sampled according to Algorithm˜1 and in Figure˜3. Additionally, note that converges rapidly to as the number of samples grows, allowing us to obtain a rigorous statistical guarantee on our stability scores as a function of the number of samples.
@@
Let for any and . For any entailment model and reasoning chain , define as in Equation˜7. Then, with probability at least , this estimate has error for all .
@@
Recall the connection between entailment stability and error detection: the lower a claim’s entailment stability , the greater its error. Consider a simple thresholding mechanism: if some estimate falls below a prescribed error threshold, then we mark the derived claim as erroneous. In the following, we demonstrate the empirical effectiveness of this procedure.
@@ 4 Evaluating ARES for Estimating Probabilistic Soundness
| Claim |
| Context Rules: H3 -> AZ; SG -> C6; C6 -> GM; VD -> H3; G8 -> VD; D8 -> U8; U8 -> DG; DG -> G8. Fact: I have D8. … | |||||||||||||||||||||||||||
| Claim 5: I use rule (VD -> H3) to derive H3 | 0.79 | 1.00 | 0.00 | 1.00 | 0.00 | 1.00 | 0.00 | 1.00 | |||||||||||||||||||
| Claim 6: I use rule (H3 -> AZ) to derive AZ | 0.82 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | |||||||||||||||||||
| Claim 7: I use rule (AZ -> SG) to derive SG | 0.00 | 0.00 | 0.00 | 1.00 | 0.00 | 1.00 | 0.00 | 0.00 | |||||||||||||||||||
| Claim 8: I use rule (SG -> C6) to derive C6 | 0.00 | 1.00 | 0.00 | 1.00 | 0.00 | 1.00 | 0.00 | 1.00 | |||||||||||||||||||
@@
For LLMs, we use GPT-4o-mini (OpenAI, 2024) and Qwen3-4B (Yang et al., 2025). For a PRM, we used Qwen2.5-Math-PRM-7B (Zhang et al., 2025). We evaluate on four datasets: PRMBench (Song et al., 2025), DeltaBench (He et al., 2025), ClaimTrees (our synthetic data), and CaptainCookRecipes (graph-based recipe dataset derived from CaptainCook4D (Peddi et al., 2024)). We evaluate using Macro-recall, Macro-precision and Macro-F1 following the literature (He et al., 2025). To compute the error threshold for the entailment scores, we first sweep over all the values that occur in the training split and select the one that maximizes Macro-F1. We repeat this process in a 5-fold cross-validation where each time we use one fold for validation and four folds for testing and report the average and standard deviation. Additional details and analyses can be found in Appendix˜C.
@@ 4.1 RQ1: Does ARES work better than baseline methods on Benchmarks?
We measure ARES’s ability to identify errors in natural reasoning chains using PRMBench and DeltaBench. With GPT-4o-mini backbone entailment model, we find that ARES achieves the best Macro-F1 scores on both datasets, shown in Table˜1. LLM-Judge performs poorly on DeltaBench while Entail-Base underperforms on PRMBench. DeltaBench’s long reasoning chains appear to confuse LLM-Judge when making holistic judgments. For Qwen3-4B, Entail-Base performs slightly better, while all other methods lag behind. Our inspection reveals that Qwen3-4B-based entailment models frequently classify next claims as entailed, suggesting limited capability for judging complex reasoning. Additional experiments in Section˜C.9 show that ARES can also achieve further improvements on top of a PRM backbone.
@@ 4.2 RQ2: In what setting does ARES identify more errors than baselines?
CaptainCookRecipes: A graph-rule-based dataset adapted from the cooking task graphs in CaptainCook4D (Peddi et al., 2024).
We designed these datasets specifically to test the core reasoning capabilities of each method, controlling for confounding variables. For example, ClaimTrees uses abstract symbols and shuffled rules to mitigate ordering bias. In both datasets, we represent the underlying rules (e.g., logical rules, recipe actions) as base claims. We then intentionally remove a key base claim—like a rule in a proof or an ingredient in a recipe—to create unsound derivations, allowing us to precisely track how the initial error propagates through the reasoning chain. Further details can be found in Appendix C.5.
Experiments on these controlled datasets confirm that ARES excels at identifying propagated errors, especially in long chains. As demonstrated in Figure 4, ARES maintains a high Macro-F1 score even as chains become very long, whereas the performance of all baseline methods deteriorates sharply after just a few steps. For example, ARES sustains a Macro-F1 score of at least 89% on chains up to 50 steps long, while other methods fall into the 30–40% range. The results in Table 1 further highlight this robust performance across our synthetic datasets, with an example shown in Table 2. We further discuss in Appendix˜B that only ARES satisfies all important desiderata for error detection while other methods fail to.
@@
4.3 RQ3: Is ARES computationally efficient?
While ARES samples multiple combinations of previous claims to check soundness, it is implemented efficiently to avoid redundant LLM calls for the same premise-hypothesis pairs. Figure 5 shows the theoretical versus actual samples used for each derived claim and complete reasoning chain, respectively. For ClaimTrees, the average total samples per example increases with chain length. With shorter chains (ClaimTrees-5), we achieve extreme efficiency at only 0.03x of theoretical samples needed. DeltaBench uses the most samples but still achieves 0.31x of theoretical samples needed. DeltaBench needing more samples indicates greater uncertainty in the entailment model’s outputted probability for this dataset. In ideal cases where the entailment model outputs only 1 or 0 for every derived claim, we need just one sample per claim each step.
To have a more direct analysis of efficiency, we conduct another analysis on the performance vs. sample size trade-off across all datasets on GPT-4o-mini in Table˜3. We find that in practice ARES’s performance is stable even with fewer samples, indicating potential further computational savings. On synthetic benchmarks ClaimTrees and CaptainCookRecipes, there is no significant difference for to , while more differences are shown for PRMBench and DeltaBench.